Grade 10Algebra

a 400kg bomb sitting at rest explodes into 3 pieces. a 150 kg piece moves off to the east with a velocity of150m/s. a 100 kg piece moves off with a velocity of 200 m/s. what is the velocity of third piece?

afaq
9 Years agoGrade 10

1 Answer

Dayitva Goel
9 Years ago
Use m1v1 + m2v2 + m3v3 = MV
 
m1 = 150 kg
m2 = 100 kg
m= 400 – (150 + 100) kg = 150 kg
M = 400 kg
 
v= 150 m/s
v2 = 200 m/s
v= ?
V = 0 m/s
 
150*150 + 100*200 +150*v3 = 400*0
22500 + 20000 + 150*v3 = 0
42500 + 150*v= 0
150*v3 = – 42500
v= – 42500 / 150
v3 = – 283.33 m/s
 
The negative sign indicates that the velocity is in opposite direction.